OpenStudy (anonymous):

Solve the system of equations by using the elimination method:

OpenStudy (anonymous):

\[4x ^{2}+3y ^{2}=12 \]

OpenStudy (anonymous):

\[5x ^{2}+6y ^{2}=30\]

OpenStudy (blurbendy):

first decide which variable you want to eliminate

OpenStudy (anonymous):

x

OpenStudy (blurbendy):

okay, in that case, the coefficients in front of x^2 need to cancel out completely. right now 4x^2 and 5x^2 will never cancel out, but if one coefficient was -20, and the other was +20, then you can cancel them out. how would you do this?

OpenStudy (anonymous):

multiply the top by 5 and the bottom by 4?

OpenStudy (blurbendy):

if you need a -20, you need to multiply of the equations by a negative. either -5 or -4

OpenStudy (blurbendy):

one of the equations*

OpenStudy (anonymous):

ok say I made the top -5 and i get \[-20x ^{2}-15y ^{2}=-60\] what do I do next?

OpenStudy (blurbendy):

good and the bottom should be: + 20x^2 +24y^2 = 120 so now add the equations -20x^2 - 15y^2 = 60 + + 20x^2 +24y^2 = 120

OpenStudy (blurbendy):

what do you get?

OpenStudy (anonymous):

\[9y ^{2}=60\]

OpenStudy (blurbendy):

perfect, now solve for y

OpenStudy (anonymous):

when I divide 9, it doesn't divide and if I simplify it, I get \[\frac{ 20 }{ 3 }\]

OpenStudy (blurbendy):

it might work out more nicely if we eliminate y first: -8x^2 -6y^2 = -24 + 5x^2 +6y^2 = 30 -3x^2 = 6 x^2 = -2 x = sqrt[-2] ?

OpenStudy (blurbendy):

so x = 2i

OpenStudy (blurbendy):

now, substitute 2i back into x for one of the equations and solve for y

OpenStudy (anonymous):

so wouldn't it just be no solution?